OpenStudy (anonymous):

Urgent, cos 3x = sen 3x solutions?

OpenStudy (anonymous):

I suppose thats sin 3x.. so, you could do two things ~ 1. divide on both sides by sin 3x to get the equation cot 3x = 1. then, you solve for 3x first, and divide the solution by 3 to get for x. 2. Bring the sin 3x to the LHS to get the equation cos 3x - sin 3x = 0, then multiply on both sides with (cos 3x + sin 3x) to get sin 6x =1, and similarly solve for 6x first, and divide the solution by 6 to get for x. By both these steps, we arrive at the general solution, \[x = \frac{ n \pi }{ 3 } + \frac{ \pi }{ 12 }\]

OpenStudy (anonymous):

Or divide by cos3x if unfamiliar with cot; leaving tan3x = 1; arctan(1) = 3x; 3x = pi/4; x = pi/12. Then repeating with the same frequency ekaknr shows.